3. You have learned that adding table salt, NaCl (58.44 g/mol, 2.16 g/cm3), to water (Kb  0.52oC/m) increases the temperature at which it boils. You decide to try cooking pasta faster at a higher temperature in boiling salty water. What increase in the boiling point do you expect if you add 4 tablespoons (1 tbsp  14.8 cm3) of salt to 16 oz of water (474 mL)?

4. The standard molar entropy of lead (II) bromide (PbBr2) is 161 J/mol  K. What is the entropy of 2.45 g of PbBr2?
1
Expert's answer
2020-03-02T06:57:19-0500

3)

Mass of 4 tablespoons of table salt ="4*volume*density=4*2.16*14.8 = 127.8 g"

converting this into moles we get

moles of NaCl added = 127.8/58.44 = 2.18

molality of the solution = 2.18/0.474 = 4.62m


"\Delta T_b=m*K_b \newline
\Delta T_b=4.62*0.52 = 2.40"


4) Molar mass of PbBr2 = 367 g

we have to find out the entropy of 2.45g of PbBr2

which is equal to "\dfrac{2.45}{367}*161 = 1.07J/molK"